OpenStudy (anonymous):

find the y-intercepts for the parabola defined by the equation y=-x^2-7x-6

jigglypuff314 (jigglypuff314):

y-intercept is when x = 0 so plug in 0 for x for y=-x^2-7x-6 you would get y = -(0)^2 - 7(0) - 6 =

OpenStudy (anonymous):

So would the answer be (0,-6) ?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

Thank you!!! :)